OpenStudy (anonymous):

Find all solutions in the interval [0, 2π). 4 sin2x - 4 sin x + 1 = 0

OpenStudy (anonymous):

@thomaster @dan815 @nincompoop @Preetha

OpenStudy (anonymous):

Solve for x on 0 ≤ x ≤2π ; 4sin² x – 4sin x +1 = 0 (2sin x – 1)² = 0 (2sin x – 1) = 0 2sin x = 1 sin x = 1/2 x = sinˉ¹(1/2) = {π/6, 5π/6} radians = {30º, 150º}

OpenStudy (anonymous):

Solve for x on 0 ≤ x ≤2π ; 4sin² x – 4sin x +1 = 0 (2sin x – 1)² = 0 (2sin x – 1) = 0 2sin x = 1 sin x = 1/2 x = sinˉ¹(1/2) = {π/6, 5π/6} radians = {30º, 150º}

OpenStudy (midhun.madhu1987):